35.3 h is what percent of 125 h?
Round appropriately.

Answer by stanbon(75887)

35.3 h is what percent of 125 h?
----
(35.3*h)/(125*h) = 35.3/125 = 0.2824 = 28.24%
--------------
Cheers,
Stan H.
---------------